OpenStudy (anonymous):

5. What are the solutions to the equation x2 + 2 = 102?

OpenStudy (pfenn1):

Solve your equation for x. Do you know how to do that?

OpenStudy (anonymous):

no

OpenStudy (pfenn1):

\[x^2 +2 =102\]Subtract 2 from each side and what do you get?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

100

OpenStudy (pfenn1):

Okay.\[x^2+2-2=102-2\]

OpenStudy (pfenn1):

\[x^2=100\] what is x?

OpenStudy (anonymous):

10

OpenStudy (pfenn1):

Great!